OpenStudy (anonymous):

Given the following function, (a) find the domain, (b) determine the vertical asymptote. (c) submit the graph of the function Please be sure to show all of your work. y = log(-x)

OpenStudy (mathteacher1729):

Do you know how to use geogebra?

OpenStudy (anonymous):

you cannot take the log of anything less than or equal to zero, so set \[-x>0\] and solve for x in one step to get \[x<0\] as your domain

OpenStudy (anonymous):

if you know what \[y=\log(x)\] looks like then \[\log(-x)\] looks the same only rotated around the y - axis

OpenStudy (anonymous):

mathteacher- i do not know how to use that...what is the site link? satellite73- i know the domain of f is (-infinity, 0) U ( 0, +infinity) but not sure what the range is and what the other one is for f^-1

OpenStudy (anonymous):

hold the phone. if the function is \[f(x)=\log(-x)\]the domain is not what you wrote, it is \[x<0\]

OpenStudy (anonymous):

the domain of log is positive numbers, and if \[-x>0\] that means \[x<0\] for sure

OpenStudy (mathteacher1729):

Hi Schlowers: download geogebra here: http://www.geogebra.org/cms/en/installers It's really easy to use. Just type your function in the input bar, and press enter.

OpenStudy (mathteacher1729):

Also, generally speaking, DOMAIN = "All the numbers you can put INTO a function." RANGE = "all the numbers which come OUT OF a function." Three things you CANNOT put in a function: 1) log of anything less than or equal to zero. 2) square root of a negative number 3) zero in the denominator Hope this helps! :D

OpenStudy (anonymous):

well that was wrong let me try again. if function is \[f(x)=\log(-x)\] domain is \[x<0\] range is \[\mathbb R\] and inverse depends on the base of the log. if base it ten inverse is \[f^{-1}(x)=-10^x\] domain is \[\mathbb R\] range is \[y<0\]

OpenStudy (anonymous):

if base is "e" then inverse is \[-e^x\] and still domain is \[\mathbb R\] range is \[(-\infty,0)\]

OpenStudy (anonymous):

ok so here is what i got: let me know if I am right or wrong. (a) x < 0 or (-infinity, 0) (b) y = log(-x) lim log (-x) x > 0 lim log (-x) x > 0^- (c) ??

OpenStudy (anonymous):

satellite- can you tell me if my answer is correct and help me again?